Suppose that you need $350,000 in 8 years. What is the amount equivalent to that sum of money today? Assume interest rate is 2% per year.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

350,000=P*(1.02)^8

P=350,000/(1.02)^8

=$298,721.63(Approx).